A cyclist increases her speed uniformly from 3m/s over a distance of 40m. Calculated the magnitude of the acceleration of the cyclist over this distance
1
Expert's answer
2020-08-12T16:44:43-0400

Condition:

"v_{0}=0 \frac ms"  

"v=3 \frac ms"

"l=40 m"

"a- ?"

Solution:

For uniformly accelerated motion:

"l=\frac{(v^2-v_0^2)}{2a}"

but "v_{0}=0 \frac ms" , then

"l=\frac {v^2}{2a}"

where

"a=\frac{v^2}{2l}"

Substitute the numbers.

"a=\frac{3^2}{2\times40}=\frac {9}{80}=0.1125\frac {m}{s^2}"

Answer: "0.1125 \frac{m}{s^2}" .